The manager of the Petroco Service Station wants to forecast the demand for unleaded gasoline next month so that the proper number of gallons can be ordered from the distributor. The owner has accumulated the following data on demand for unleaded gasoline from sales during the past 10 months:

Months

Demand

October

800

November

725

December

630

January

500

February

645

March

690

April

730

May

810

June

1200

July

980

Compute the following:

a- MA3

b- MA4

c- WMA3 (0.5, 0.3, 0.2)

d- Exponential Smoothing forecast (α=0.3)

e- Adjusted Exponential Smoothing forecast (α=0.3, β=0.2)

Part 2

Using MAD, determine which of these five forecasts is more accurate.
Answer #1

a)

Months Demand Forecast Error abs(error)
October 800
November 725
December 630
January 500 718.3333 -218.333 218.3333
February 645 618.3333 26.66667 26.66667
March 690 591.6667 98.33333 98.33333
April 730 611.6667 118.3333 118.3333
May 810 688.3333 121.6667 121.6667
June 1200 743.3333 456.6667 456.6667
July 980 913.3333 66.66667 66.66667
Total 1106.667
MAD 158.0952

b.

c.

d.

e.

F(t) = FIT(t-1) + alpha*(A(t-1) - FIT(t-1) )

T(t) = T(t-1) + beta * (F(t) - FIT(t-1) )

FIT(t) = F(t) +T(t)

Part - 2

Using MAD c- WMA3 (0.5, 0.3, 0.2) is best as it has lowest MAD